OpenStudy (precal):

Let F=f+g and G=3f-g find F'(1)

OpenStudy (akshay_budhkar):

Do u have f and g given?

OpenStudy (precal):

I have a graph, would you like me to upload it?

OpenStudy (akshay_budhkar):

yea sure i will have a look at it

OpenStudy (precal):

Ok, give me a minute I have to scan it.

OpenStudy (anonymous):

Akshay had a good eyes :P

OpenStudy (precal):

Thanks for waiting. I think F ' = f ' + g ' G ' = 3f ' - g' not sure about G '

OpenStudy (precal):

I am not sure what to do with F=f+g

OpenStudy (anonymous):

This is the hard way: Use interpolation to find the value of f(x) and g(x) in polynomial form and then use the derivatives.

OpenStudy (precal):

sorry not sure how to do that. Interpolation?

OpenStudy (akshay_budhkar):

@Fool o_O??? What is the easy way? I would go for finding f and g as a general function ( it is a line) and add the general equations then differentiate and the substitute 1

OpenStudy (precal):

ok let me write the equations

OpenStudy (precal):

so, I only need the equation where x=1 so f(x)=-3x+10 now let me do g(x)

OpenStudy (precal):

g(x) is x+1 F=f+g F=-3x+10+x+1 F=-2x+11 F'=-2 F'(1) = -2 correct?

OpenStudy (precal):

yes, I am right. Lookedin the back of my book. Thanks for your help

OpenStudy (akshay_budhkar):

hehe when u wrote ur function f i got it u will get your answer!! i was sure of it =D